OpenStudy (anonymous):

HELP PLEASE! Verify the identity: sin(x+pi/2)=cos (x)

OpenStudy (anonymous):

\[\sin \left(x + \frac{ \pi }{ 2 } \right) = \left( \sin x \right)\left( \cos \frac{ \pi }{ 2 } \right) + \left( \cos x \right)\left( \sin \frac{ \pi }{ 2 } \right)\]\[= \left( \sin x \right)(0) + (\cos x)(1)\]

OpenStudy (anonymous):

lhs= sin(x+pi/2) =SINX COS PI/2+COS X SIN PI/2 =SIN X *0 + COS X*1 =COS X =RHS Hence proved

OpenStudy (anonymous):

could you help with another one too?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok tcarroll

OpenStudy (anonymous):

Post your question

OpenStudy (anonymous):

thanks! See you in the new post!